Grade 12Analytical Geometry

find the electic field stregth due to a charged circular arc at its centre

yuvraj
9 Years agoGrade 12

1 Answer

Pranav
9 Years ago
Consider an element on the ring(having charge per unit length l) having charge dq.Now dq=l*r *da.(da is angle surrender by the element at centre and r is the radius)let de be electric field at centre due to this element de=kdq/r^2=klrda/r^2=klda/rIntegrate this with limit 0 to b.where b is the angle suspended by whole arc.The result will be E=2klsin (b/2)/r